OpenStudy (anonymous):

Polynomial Functions Use your graphing calculator to graph the following functions. Using the definition of Polynomial Functions, determine which of the functions are polynomials. 1. y=x^2-3x+1 2. y=|x| 3. y=sqrt(x) 4. y=5 5. y=(x-2)(x-1)(x+5) 6. y=3/(x-2) 7. y=(x+2)/4 a. Polynomial Function b. Not a Polynomial Function

OpenStudy (anonymous):

can u unblock me anytime soon

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

y=x^2-3x+1 quadratic polynomial

OpenStudy (anonymous):

y=5 constant polynomial

OpenStudy (anonymous):

y=(x-2)(x-1)(x+5) cubic polynomial

OpenStudy (anonymous):

y=(x+2)/4 linear polynomial

OpenStudy (anonymous):

the rest are not

OpenStudy (anonymous):

whats the difference between 6 and 7? why is one polynomial and one isn't?